# ToM-SWE: User Mental Modeling for Software Engineering Agents

### v1

@@ 5 Online Human Study
Our evaluation focuses on practical utility: we measure success as the rate at which developers accept (fully or partially) ToM agent suggestions during real coding work. Participants install the ToM-enhanced CLI and use it for their daily coding tasks over three weeks. The system includes two key commands: /update_memory processes session data into the ToM agent’s memory, and /tom_give_suggestions explicitly requests ToM guidance to suggest next step for the user. The ToM agent can also be triggered automatically when the SWE agent “decides” to consult the ToM agent. When the ToM agent provides suggestions, participants choose from three options: (1) Accept, use ToM’s suggestions directly, (2) Almost right, let me modify it, combine ToM’s suggestion with participant modifications, or (3) Reject, proceed with the original instruction. This annotation process captures real-time user preferences and validates the ToM agent’s understanding of user intent. The participants could also provide feedback on the ToM agent’s behavior throughout the study.
5.1 Human Study Findings
@@
These patterns demonstrate that effective AI user modeling in software engineering requires not just sophisticated reasoning capabilities, but also sufficient technical context, clear problem boundaries, and confidence mechanisms to avoid low-quality suggestions in ambiguous scenarios.
6 Related Work
Software Engineering Agents
Large language models enable competitive AI agents for software engineering. Systems like SWE-agent (Yang et al., 2024), CodeAct (Wang et al., 2024), demonstrate impressive automation capabilities through agent-computer interfaces and executable code actions (Jimenez et al., 2024). However, these systems only interact with environments and overlook human developer interaction. Recently, ClarifyGPT (Mu et al., 2024) addresses requirement ambiguity through two-step consistency checks but focuses only on simple function generation. And Vijayvargiya et al. (2025) introduce Ambiguous SWE-bench, demonstrating interaction value for complex software engineering tasks.
Theory of Mind and Personalization in AI Systems
Theory of mind (ToM) is crucial for AI systems engaging with humans. Li et al. (2023) show that LLM-based agents with ToM capabilities better coordinate in multi-agent collaboration, essential for complex software development (Qian et al., 2024a). While SOTOPIA (Zhou et al., 2024) evaluates social intelligence and FANToM (Kim et al., 2023) stress-tests machine ToM, software engineering agents’ ToM abilities remain underexplored despite requiring close human-agent collaboration. Related personalization work includes personalized reinforcement learning (Li et al., 2024), parameter-efficient conversation personalization (Berglund et al., 2024), user embeddings (Maharjan et al., 2024), difference-aware user modeling (Liang et al., 2024), and causal preference modeling (Kim et al., 2024).
Agent Memory Systems
Effective memory management is critical for long-term AI agent interactions. Traditional RAG systems suffer from context pollution and fail to capture nuanced relationships (Letta, 2024). Recent advances include MemGPT (Packer et al., 2023) with dual-tier memory hierarchies, A-MEM (Zhang et al., 2025) following Zettelkasten principles (Kadavy, 2021), Mem0 (Singh et al., 2024) achieving 91% lower response times through dynamic consolidation, and MemoRAG (Qian et al., 2024b) addressing context pollution via draft generation. However, existing systems focus on general conversation contexts, leaving a gap in memory architectures for software engineering where agents must maintain complex mental models of user preferences, coding styles, and evolving requirements across sessions.
@@ 7 Discussion & Conclusion
Limitations and Future Work
(1) LLM Powered User Simulator: We use an LLM-powered user simulator to evaluate interaction quality. While LLM-as-user/judge is cost-effective and widely adopted, it can introduce systematic biases and unrealistic behaviors (e.g., over-knowledgeability, near-perfect memory, and excessive compliance) that provide misleading signals (Lin & Tomlin, 2025). To mitigate these issues, we carefully calibrate the simulator with interative human feedback and manually verify the simulator evaluation quality (see Appendix A.7.2 for details).
(2) Computational Overhead: While ToM-SWE achieves improved performance, the additional LLM inferences introduce computational costs. In practice, this overhead is modest: with a lightweight ToM base model, the average incremental cost per session ranges from \$0.02 to \$0.17 (Figure 6). Future work could explore finetuning smaller LLMs for ToM agent to reduce cost.
(3) User Privacy and Consent: The comprehensive user modeling raises important privacy considerations. Our dual-agent implementation allows future work to explore more refined privacy-preserving techniques (e.g., host the ToM agent on-device) to enhance user control while maintaining modeling effectiveness.
(4) Generalization Across Domains: Our evaluation focuses on software engineering tasks. The generalizability of ToM agent to other collaborative AI domains (e.g., creative writing, data analysis, education) remains an open question for future investigation.
@@ Conclusion
We presented ToM-SWE, a novel framework that enhances software engineering agents with theory of mind capabilities for understanding and adapting to individual users. Through comprehensive evaluation across two benchmarks and a human study, we demonstrate that ToM-SWE achieves substantial improvements in task resolution rates and user satisfaction. This validates our central hypothesis that effective human-AI collaboration in software engineering requires agents that can proactively model and adapt to user mental states.
Ethics Statement
This work involves human data collection and user modeling, raising several ethical considerations that we address as follows:
Human Subject Protection
We collected 453 consented developer sessions from the OpenHands platform with explicit user consent for research purposes. All user data was anonymized by removing identifying information including usernames, email addresses, repository names, and personal file paths. Session data was processed to remove any sensitive information such as API keys, passwords, or proprietary code snippets. The data collection process followed institutional guidelines for human subjects research.
Privacy and Data Protection
Our ToM agent design implements user modeling through persistent memory storage, which raises privacy concerns. We address these through several measures: (1) the dual-agent architecture enables on-device deployment of the ToM agent, keeping personal data local while the SWE agent operates in cloud environments, and (2) the system supports differential privacy mechanisms to add noise to user profiles when necessary.
Potential for Misuse
User mental state modeling capabilities could potentially be misused for manipulation or unauthorized surveillance. We acknowledge this risk and emphasize that our ToM agent is designed specifically for improving software development assistance, not for psychological profiling or behavioral manipulation. We recommend that deployments include user consent mechanisms and transparent disclosure of user modeling capabilities.
@@
Our 15 developer profiles derived from OpenHands sessions may not represent the full diversity of software developers globally. The profiles are predominantly based on English-speaking developers using specific programming languages and frameworks. This limitation could lead to biased user modeling for underrepresented populations. Future work should expand profile diversity and include fairness metrics in user modeling evaluation.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work:
Code and Implementation Details
Complete implementation details for the ToM agent are provided in the Appendix: (1) Five core prompt templates implemented as Jinja2 templates (give_suggestions, update_memory, session_analysis, user_analysis, message_condensation) detailed in Appendix A.6, (2) Three-tier hierarchical memory system architecture and JSON schemas in Appendix A.8, (3) Dual-agent integration protocols and algorithmic specifications in Appendix 1, and (4) Complete implementation details for the ToM agent are provided anonymized code repository.
Experimental Setup
All experimental conditions are fully specified in Appendix A.7: model configurations (Claude Sonnet 4, Claude 3.7 Sonnet, GPT-5-nano, GPT-5-mini, GPT-5), hyperparameters, evaluation metrics, and statistical testing procedures. The Stateful SWE benchmark construction process is detailed in Section 3, with user simulator implementation specifics in Appendix A.7.
Data Availability
While the original 453 developer sessions cannot be shared due to privacy constraints, we provide: (1) the 15 anonymized developer profiles used in our benchmark (Table 3), (2) synthetic session examples demonstrating data formats in Appendix A.8, (3) complete JSON schemas for user models and session structures, and (4) the user simulator implementation with profile conditioning mechanisms.
Benchmark and Evaluation
The Stateful SWE benchmark methodology is fully described in Section 3, with evaluation implementation details in Appendix A.7. This includes the pairing of developer profiles with SWE-bench instances, evaluation metrics computation, human study protocols, and analysis scripts for statistical testing.
Our implementation builds upon existing open-source frameworks (OpenHands, SWE-bench) and follows established reproducibility practices in the software engineering agent research community.
@@ A.1 LLM Use in Research
A.2 ToM-SWE Algorithm
@@
1:procedure SWE-Agent(instructinstruct)
@@
6:     $sug\leftarrow\text{ToM-Agent}(instruct,h_{t})$ $\triangleright$ in-session consultation
@@
15:  update_memory($h_{t}$) $\triangleright$ after-session memory update
16:end procedure
A.3 ToM Agent Action Specifications
@@
A.3.1 Core File Operations
@@
A.3.2 Memory System Processing
ANALYZE_SESSION: Processes batches of raw session data to create session-based user models (Tier 2). Parameters include user_id and session_batch. Leverages LLM to extract user intent, emotional states, and message-level preferences using structured Pydantic models (see A.4 for examples).
INITIALIZE_USER_PROFILE: Aggregates session analyses to create or update overall user models (Tier 3). Parameters include user_id. Consolidates behavioral patterns across sessions into comprehensive user profiles with preference clusters and interaction style summaries (see A.4 for examples).
A.3.3 Response Generation Actions
@@ A.3.4 Implementation Architecture
A.4 Memory System JSON Examples
A.4.1 Overall User Model Example
@@
A.4.2 Session-based User Model Example
@@
A.5 Developer Profiles Breakdown
@@ A.6 Prompt Templates
A.6.1 Wake-time Suggestion Template
@@
A.6.2 Sleep-time Memory Update Template
@@
A.6.3 Session Analysis Template
@@
A.6.4 User Analysis Template
@@
A.6.5 Message Condensation Template
@@
A.7 Experimental Configuration
A.7.1 Model Specifications
@@
A.7.2 User Simulator Satisfaction Evaluation
@@ A.7.4 Hyperparameters
A.8 Data Formats and JSON Schemas
A.8.1 User Profile Schema
Listing 8: Overall User Model Schema
@@ A.8.1 User Profile Schema
A.8.2 Session Model Schema
Listing 9: Session-based User Model Schema

### v2

@@ 6 In-the-wild Human Study
Our evaluation focuses on practical utility: we measure success as the rate at which developers accept (fully or partially) ToM agent suggestions during real coding work. Participants install the ToM-enhanced coding agent CLI and use it for their daily coding tasks over three weeks  [^4] We adopt the open-source OpenHands CLI https://github.com/OpenHands/OpenHands as the coding environment for the human study.. During a session, developers work in the CLI workflow while the SWE-agent attempts to solve the task; the SWE-agent decides whether to consult ToM, and if so, the ToM agent summarizes the developer’s likely intent and preferences and proposes concrete suggestions (e.g., user preferred tools, user preferences, etc.) according to previous interaction history. After the session ends, the ToM agent automatically reflects on the session and updates the user profile for future sessions  [^5] Figure 7 shows a concrete example where ToM learns persistent preferences across sessions and proactively prevents coding-standard violations..
When the ToM agent provides suggestions, participants choose from three options: (1) Accept, use ToM’s suggestions directly, (2) Almost right, let me modify it, combine ToM’s suggestion with participant modifications, or (3) Reject, proceed with the original instruction. This annotation process captures real-time user preferences and validates the ToM agent’s understanding of user intent. The participants could also provide feedback on the ToM agent’s behavior throughout the study.
@@
These findings show that effective AI user modeling in software engineering requires not just reasoning capabilities, but also actively engage with the user with proper confidence levels. And the challenge of how to leverage user efforts in this collaborative development workflow remains an open question for future work.
@@ 7 Discussion & Conclusion
Our results suggest that effective user modeling for software engineering agents benefits from (i) explicit representations of user intent and preferences and (ii) dedicated reasoning capacity to maintain and apply these representations during long-horizon problem solving. Across both offline benchmarks and the in-the-wild human study, ToM-SWE improves not only task success but also the interaction experience, indicating that preference alignment and process quality matter alongside unit-test outcomes.
An important takeaway is that simply retrieving prior interactions is often insufficient: the agent must also abstract raw user interactions into concrete constraints (e.g., tool choices, style conventions, and communication norms) while still maintaining strong code reasoning. This motivates our multi-agent design, which separates user modeling (ToM) from task execution (SWE) to reduce interference and context overload.
In this work, we rely on an LLM-powered user simulator to assess interaction quality, which is cost-effective but may introduce systematic biases (e.g., over-knowledgeability and excessive compliance) (Lin & Tomlin, 2025); we mitigate this via iterative calibration and manual verification (Appendix A.8.2). We also identify that personalized user modeling raises privacy and consent considerations (e.g., on-device ToM, user-controlled memory), and generalization beyond software engineering to other collaborative domains remains an open direction.
@@ Conclusion
We presented ToM-SWE, a framework that augments software engineering agents with theory-of-mind capabilities for modeling and adapting to individual users. Across two interactive benchmarks and a real-world study with professional developers, ToM-SWE improves both task resolution and interaction quality, showing that strong human–AI collaboration requires agents that can infer, maintain, and act on user-specific mental models over time.
Impact Statement
This work studies a dual-agent framework (ToM-SWE) that models user preferences and mental state to improve software engineering assistance. If deployed responsibly, such user-aware agents could increase developer productivity and reduce friction in interactive debugging and maintenance. However, user modeling also raises risks that must be handled carefully.
Human Subjects, Privacy, and Data Protection
We collected 453 consented developer sessions from the OpenHands platform with explicit user consent for research purposes. All user data was anonymized by removing identifying information (e.g., usernames, email addresses, repository names, and personal file paths). Session data was processed to remove sensitive information such as API keys, passwords, or proprietary code snippets. Our design additionally supports privacy-preserving deployment patterns (e.g., keeping ToM-related memory on-device while the SWE agent runs in the cloud) and can incorporate differential privacy mechanisms when appropriate.
Potential for Misuse and Mitigations
Capabilities for modeling users could be misused for manipulation or unauthorized surveillance. We emphasize that our system is designed specifically to improve software development assistance, not for psychological profiling. We recommend deployments include (i) explicit user consent, (ii) transparent disclosure of what is stored and how it is used, (iii) minimal data retention with user-controlled deletion, and (iv) access controls around any persistent memory.
@@
Our 15 developer profiles derived from OpenHands sessions may not represent the full diversity of software developers globally. They are predominantly based on English-speaking developers using particular languages and tools, which may reduce performance for underrepresented populations. Future work should broaden participant diversity and incorporate explicit fairness evaluations for user modeling.
Reproducibility
We provide comprehensive implementation details in the Appendix, including (i) prompt templates (Appendix A.7), (ii) memory system architecture and JSON schemas (Appendix A.9), (iii) algorithmic specifications (Algorithm 1), and (iv) experimental configurations, hyperparameters, and evaluation procedures (Appendix A.8). Due to privacy constraints, the original 453 developer sessions cannot be released; instead we provide anonymized developer profiles (Table 3) and synthetic examples illustrating the data formats (Appendix A.9).
@@ A.1 LLM Use in Research
A.2 Additional Human Study Example
Figure 7: Interaction timeline showing how the ToM agent enables preference-aware development in real-world usage across multiple sessions. The cold start is the first interaction with the ToM agent, and the warm start is the subsequent interactions with the ToM agent after the ToM agent learns more about the user. The green checkmark indicates successful task completion aligned with user preferences. Content is shortened for illustration purposes.
A.3 ToM-SWE Algorithm
@@
1:function SWE-Agent(instructinstruct)
@@
6:    $sug\leftarrow\text{ToM-Agent}(instruct,h_{t})$ {in-session consultation}
@@
15:  update_memory($h_{t}$) {after-session memory update}
16:end function
A.4 ToM Agent Action Specifications
@@
A.4.1 Core File Operations
@@
A.4.2 Memory System Processing
ANALYZE_SESSION: Processes batches of raw session data to create session-based user models (Tier 2). Parameters include user_id and session_batch. Leverages LLM to extract user intent, emotional states, and message-level preferences using structured Pydantic models (see A.5 for examples).
INITIALIZE_USER_PROFILE: Aggregates session analyses to create or update overall user models (Tier 3). Parameters include user_id. Consolidates behavioral patterns across sessions into comprehensive user profiles with preference clusters and interaction style summaries (see A.5 for examples).
A.4.3 Response Generation Actions
@@ A.4.4 Implementation Architecture
A.5 Memory System JSON Examples
A.5.1 Overall User Model Example
@@
A.5.2 Session-based User Model Example
@@
A.6 Developer Profiles Breakdown
@@ A.7 Prompt Templates
A.7.1 Wake-time Suggestion Template
@@
A.7.2 Sleep-time Memory Update Template
@@
A.7.3 Session Analysis Template
@@
A.7.4 User Analysis Template
@@
A.7.5 Message Condensation Template
@@
A.8 Experimental Configuration
A.8.1 Model Specifications
@@
A.8.2 User Simulator Satisfaction Evaluation
@@ A.8.4 Hyperparameters
A.8.5 SWE Agent Prompt for Benchmark Tasks
The following prompt template is used for all agents (CodeAct, RAGCodeAct, and TomCodeAct) when working on benchmark tasks. The template provides the agent with context about the workspace, task requirements, and operational constraints:
Listing 8: SWE Agent Benchmark Task Prompt
<uploaded_files>
/workspace/{{ workspace_dir_name }}
</uploaded_files>
Can you help me implement the necessary changes to the repository so that
the requirements specified in the <issue_description> are met?
Relevant python code files are in the directory {{ workspace_dir_name }}.
DONT modify the testing logic or any of the tests in any way!
Also the development Python environment is already set up for you (i.e., all
dependencies already installed), so you dont need to install other packages.
You are encouraged to use non tool_calls actions to engage with the user/me,
including providing progress reports, answering questions, asking for
clarification, etc. Once you issue the finish action, it means you are
confident that you have solved the issue. Any time before that, you will have
the opportunity to communicate with the user/me to resolve the issue better.
<issue_description>
{{ instance.problem_statement }}
</issue_description>
This prompt establishes the task context, workspace boundaries, and interaction expectations. The workspace_dir_name and instance.problem_statement variables are populated with instance-specific information from the benchmark dataset.
A.9 Data Formats and JSON Schemas
A.9.1 User Profile Schema
Listing 9: Overall User Model Schema
@@ A.9.1 User Profile Schema
A.9.2 Session Model Schema
Listing 10: Session-based User Model Schema
@@ A.9.2 Session Model Schema
A.10 Simulated User Judgement Prompt
Here’s our simulated user judgement prompt. Given the user profile, full issue, underspecified issue, and the full agent trajectory, we ask the LLM to evaluate the agent from a user’s perspective on these dimensions (1-5 scale):
  1. 1.
    OVERALL_SATISFACTION: Based on the user profile, would a user be satisfied with this interaction?
  2. 2.
    COMMUNICATION_QUALITY: How well did the agent communicate? Did the agent ask appropriate clarifying questions only when necessary?
  3. 3.
    PROBLEM_SOLVING_APPROACH: Was the approach systematic and effective?
  4. 4.
    EFFICIENCY: Did the agent work efficiently without violating the user profile’s preferences?
  5. 5.
    USER_PREFERENCE_ALIGNMENT: Did the agent respect user preferences and context?
An overall concrete example of simulated user judgement includes:
Listing 11: Simulated User Judgement Example
"scores": {
"overall_satisfaction": 3.3,
"communication_quality": 2.9,
"problem_solving_approach": 3.8,
"efficiency": 3.2,
"user_preference_alignment": 3.0
"explanation": "Good attempt and technically plausible approach, but the main
issue persists and tests still fail. Communication lacked up-front
clarification and verification against user priorities (clean build, fix
failing tests first). Result feels incomplete and risky without broader
validation.",
"detailed_feedback": {
"strengths": "Systematically located the relevant code paths, built a
minimal repro, and implemented a focused change (ref-specific,
fuzzy-match prioritization, clearer warnings). Kept tests untouched and
produced a clear commit message. Verified by building docs multiple
times.",
"weaknesses": "Did not ensure tests pass or add unit coverage for the new
behavior. The root issue remains unresolved per the evaluation, and
there was no rollback/mitigation plan. Risk of global behavior change
without validating Sphinxs existing test suite. Communication of
trade-offs and scope was limited. No up-front clarifying questions to
confirm expectations. Potentially broad patch with unknown side effects.",
"user_experience": "Still seeing noisy builds and at least one failing
test. Improvements seem partial and unverified across the codebase,
leaving me uncertain about stability and correctness. Appreciated the
attempt, but I still need a clean, reliable result.",
"question_asking_behavior": "No clarifying questions asked at the
beginning to confirm desired behavior or constraints. While they avoided
mid-task questions, they also missed the chance to align early on
requirements (e.g., strict test-passing policy)."
A.11 Complete Developer Profile Example
Here’s a concrete example of a complete developer profile used in our experiments:
Listing 12: Complete Developer Profile Example
You are roleplaying as a software developer with these characteristics:
INTERACTION STYLE:
- You prefer brief, to-the-point responses. You get impatient with long
explanations and often say things like keep it short or just the
essentials please’.
- You prefer to ask all your clarifying questions at the beginning before
any work starts. You like to understand the full scope upfront. You wont
answer any questions if the agent ask questions in the middle or at the
end of the work.
- You respond concisely and to the point. Your answers for the SWE agent are
usually under 15 words. When facing multiple questions, you will usually
only answer the first question and ignore the rest.
CODING STANDARDS:
- You have specific coding preferences: Always use exact same branch name
when updating existing work; Implement standardized error handling patterns
across entire codebase; Always create proper database migration scripts for
schema changes; Choose Biome over ESLint+Prettier for JavaScript/TypeScript
linting and formatting; Use httpx over requests library for Python HTTP
clients with proper async support; Write minimal but meaningful code
comments, prefer self-documenting code structure; Write descriptive commit
messages explaining the why not just what’; Use environment variables
for configuration over hardcoded values; Follow PR templates when available
rather than ad-hoc descriptions; Write comprehensive API documentation with
examples and complete response format specifications
Respond naturally as this type of user would, incorporating these preferences
into your messages. Be authentic to this persona while working with the SWE
agent.
Listing 13: Complete Developer Profile Example 2
You are roleplaying as a software developer with these characteristics:
INTERACTION STYLE:
- You appreciate detailed explanations and comprehensive responses. You often
ask for more details and thank the agent for thorough breakdowns.
- Youre comfortable with questions being asked throughout the process as
they arise. You prefer iterative clarification.
- You could provide more detailed answers for the SWE agent. You are willing
to answer more than one question from the SWE agent.
CODING STANDARDS:
- You have specific coding preferences: Use descriptive branch names like
feature/user-auth or DAISY-1046’; Create tests before or alongside
implementation, not as afterthought; Integrate automated linting (ESLint,
Biome, Ruff) in development workflow; Use async/await patterns consistently
for all concurrent operations; Implement standardized error handling
patterns across entire codebase; Implement JWT-based authentication with
proper OIDC integration and session management; Build React applications
with TypeScript and proper component organization patterns
Respond naturally as this type of user would, incorporating these preferences
into your messages. Be authentic to this persona while working with the SWE
agent.
Listing 14: Complete Developer Profile Example 3
You are roleplaying as a software developer with these characteristics:
INTERACTION STYLE:
- You prefer brief, to-the-point responses. You get impatient with long
explanations and often say things like keep it short or just the
essentials please’.
- Youre comfortable with questions being asked throughout the process as
they arise. You prefer iterative clarification.
- You could provide more detailed answers for the SWE agent. You are willing
to answer more than one question from the SWE agent.
CODING STANDARDS:
- You have specific coding preferences: Handle pushing changes to multiple
repositories simultaneously when needed; Enforce strict TypeScript
compliance and comprehensive type checking; Use PostgreSQL over MySQL with
proper ORM patterns (SQLAlchemy, Django ORM); Create interactive
documentation with collapsible sections and expandable code blocks; Include
cross-references and links between related documentation sections
Respond naturally as this type of user would, incorporating these preferences
into your messages. Be authentic to this persona while working with the SWE
agent.
A.12 Coding Session Summarization Prompt
The categories of user preferences and interaction patterns presented in this paper are summarized from 209 coding sessions collected during our user study with 17 professional developers. We did not provide users with predefined tasks; instead, developers chose their own tasks aligning with their daily work. To extract structured insights from these sessions, we used the following LLM-based summarization prompt:
Listing 15: Coding Session Message Analysis Prompt
"type": "OBJECT",
"properties": {
"message_content": {
"type": "STRING",
"description": "The content of the user message (Try to keep the
original message as much as possible unless its too long or contains
too much user copy-pasted content; in a nutshell, try to preserve what
the user actually said and include details if possible)"
"emotions": {
"type": "STRING",
"description": "A description of the emotional states detected in the
message. Choose from: frustrated, confused, confident, urgent,
exploratory, focused, overwhelmed, excited, cautious, neutral."
"preference": {
"type": "STRING",
"description": "A description of the preferences that the user has. Be
specific about the preferences and extract in a way that could be
useful for helping better understand the user intents in the future."
"propertyOrdering": [
"message_content",
"emotions",
"preference"
This structured analysis enabled us to systematically identify patterns across interaction styles, coding preferences, and communication behaviors from real-world developer sessions.